OpenStudy (anonymous):

The population of a local species of dragonfly can be found using an infinite geometric series where a1 = 30 and the common ratio is 2/5. Write the sum in sigma notation and calculate the sum that will be the upper limit of this population.

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (anonymous):

this divergent right?

OpenStudy (kc_kennylau):

\[\Large\sum_{k=1}^\infty\left[30\times\left(\frac25\right)^{n-1}\right]\]

OpenStudy (kc_kennylau):

since |the ratio|<1, it's convergent